11. Show that if P is a prime ideal of D, then Dp = {a/b € Q(D) | b & P} is an integral domain with DCDP SQ(D).

All I know is that this problem entails localization of a prime...
Answer #1

Solution:-

Given that

So, you need to prove that, Pc D is prime

iff \frac{D}{P} is an integral domain

D_p=\frac{a}{b}

Now, firstly let suppose Pc D , & try to prove that \frac{D}{P} is an integral domain.

we have to prove commutative ving and ith multiplicative identity & non zero divisor

but D_p is already an commutativering with identity D.

D_p=\left \{ \frac{a}{b}\in Q(D) \right \},a,b\in Q(D)

i.e., a,b\in P for non zero element,

a + P & b + P be non zero elements, a,b\in P

Since P is an prime, a\cdot b\notin P

a\cdot b+ P is also an non zero as we started.

any two non zero elements of D_p have non zero product, i.e., non zero divisor.

Therefore D_p is an integral domain.

conversly; we assume D\subseteq D_p

Such that, D_p=\frac{D}{p} is an integral domain.

So, we need to prove that P is an prime ideal.

for this we need to show;

if ab\in p then neither a\in p or b\in p

but if ab\in p then ab+p=p

but we know, ab+p=(a+p)(b+p) is zero in D_p

but D_p is an integral domain by assumption

It can't have divisors

This proves that P is prime ideal for D

Hence proved.

Thanks for supporting...

Please give positive rating...